# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v4

@@ Acknowledgments
Table S5: TESSERA provides unprecedented ease of use, scale, and accuracy. Unlike other remote sensing foundation models, TESSERA provides analysis-ready and gap-free outputs (R), is open (O), and provides global (G) annual (A) coverage at 10 m resolution (10 m). It uses spectral-temporal features (ST) at pixel level (P), is invariant to clouds (CI), provides a compressed representation (CR), is oriented to land features (LO), and includes both multispectral and radar information (MS). Codes: $\checkmark$= Yes, $\times$= No, ? = Unknown, p= Potentially true, .. = N/A.
@@
GeoKR (2021)  (?) $\times$ ? $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
CVPRW (2021)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
GASSL (2021)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ .. $\checkmark$ $\times$
SeCo (2021)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
DINO-MM (2022)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\checkmark$
SatMAE (2022)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ p .. p $\times$
RS-BYOL (2022)  (?) $\times$ .. .. $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ .. $\times$ $\checkmark$
GeCo (2022)  (?) $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. p $\times$
RingMo (2022)  (?) $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
RVSA (2022)  (?) $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
RSP (2022)  (?) $\times$ .. .. $\times$ p $\times$ $\times$ p $\checkmark$ $\checkmark$ $\times$
MATTER (2022)  (?) $\times$ .. .. $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
CSPT (2022)  (?) $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
CVPRW (2022)  (?) $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
BFM (2023)  (?) $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
TOV (2023)  (?) ? .. $\times$ $\times$ $\checkmark$ $\times$ ? $\times$ ? $\checkmark$ $\times$
CMID (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
RingMo-Sense (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ ? $\checkmark$ $\checkmark$
lal-SimCLR (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
CACo (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
SatLas (2023)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
GFM (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Scale-MAE (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
DINO-MC (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
CROMA (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Cross-Scale MAE (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
DeCUR (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Presto (2023)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CtxMIM (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
FG-MAE (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Prithvi (2023)  (?) $\times$ .. $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
RingMo-Lite (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
IGARSS (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
EarthPTi (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
USat (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
AIEarth (2023)  (?) p ? $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Clay (2023)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Hydro (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
U-Barn (2023)  (?) $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
GeRSP (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
SwiMDiff (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
OFA-Net (2023)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
SML-FR (2023)  (?) p p $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Spectral-GP (2024)  (?) $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
S2MAE (2024)  (?) $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
SatMAE++ (2024)  (?) $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
msGFM (2024)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
SkySense (2024)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
MTP (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
DOFA (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
MMEarth (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
LeMeViT (2024)  (?) $\times$ .. $\times$ $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
SoftCon (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
RS-DFM (2024)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
A2MAE (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
HyperSIGM (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
SelectiveMAE (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
OmniSat (2024)  (?) $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
MM-VSE (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
MA3E (2024)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
SpectralEarth (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
SenPa-MAE (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
RingMo-Ae (2024)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
SAR-JEPA (2024)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
PIS (2024)  (?) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ .. .. .. .. .. ..
OReole-FM (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
PIEVIT (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
SatVisionTOA (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
RS-vHeat (2024)  (?) $\times$ .. ? ? ? ? $\checkmark$ ? $\checkmark$ ? $\checkmark$
Prithvi-EO-2.0 (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
AnySat (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
WildSAT (2024)  (?) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
SeaMo (2024)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
FoMo (2025)  (?) $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
SatMamba (2025)  (?) $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Galileo (2025)  (?) p p $\checkmark$ ? $\checkmark$ $\times$ ? ? $\checkmark$ $\checkmark$ $\checkmark$
SatDiFuser (2025) (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
RoMA (2025)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Panopticon (2025)  (?) $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
FedSense (2025)  (?) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Copernicus-FM (2025)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
HyperFree (2025)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
DynamicVis (2025)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
FlexiMo (2025)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Google Satellite (2025) (?) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
RingMoE (2025)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$

### v5

@@ A.9 Interactive Habitat Mapping Tool
Table 5: TESSERA provides unprecedented ease of use, scale, and accuracy. Unlike other remote sensing foundation models, TESSERA provides analysis-ready and gap-free outputs (R), is open (O), and provides global (G) annual (A) coverage at 10 m resolution (10 m). It uses spectral-temporal features (ST) at pixel level (P), is invariant to clouds (CI), provides a compressed representation (CR), is oriented to land features (LO), and includes both multispectral and radar information (MS). Codes: $\checkmark$= Yes, $\times$= No, ? = Unknown, p= Potentially true, .. = N/A.
@@
GeoKR (2021) [121] $\times$ ? $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
CVPRW (2021) [122] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
GASSL (2021) [23] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ .. $\checkmark$ $\times$
SeCo (2021) [123] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
MOSAIKS (2021) [124] $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ ?
DINO-MM (2022) [125] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\checkmark$
SatMAE (2022) [16] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ p .. p $\times$
RS-BYOL (2022) [126] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ .. $\times$ $\checkmark$
GeCo (2022) [127] $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. p $\times$
RingMo (2022) [128] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
RVSA (2022) [129] $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
RSP (2022) [14] $\times$ .. .. $\times$ p $\times$ $\times$ p $\checkmark$ $\checkmark$ $\times$
MATTER (2022) [130] $\times$ .. .. $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
CSPT (2022) [131] $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
CVPRW (2022) [132] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
BFM (2023) [133] $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
TOV (2023) [134] ? .. $\times$ $\times$ $\checkmark$ $\times$ ? $\times$ ? $\checkmark$ $\times$
CMID (2023) [135] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
RingMo-Sense (2023) [136] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ ? $\checkmark$ $\checkmark$
lal-SimCLR (2023) [137] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
CACo (2023) [138] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
SatLas (2023) [15] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
GFM (2023) [139] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Scale-MAE (2023) [140] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
DINO-MC (2023) [141] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
CROMA (2023) [142] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Cross-Scale MAE (2023) [143] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
DeCUR (2023) [144] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Presto (2023) [19] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CtxMIM (2023) [145] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
FG-MAE (2023) [146] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Prithvi (2023) [17] $\times$ .. $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
RingMo-Lite (2023) [147] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
IGARSS (2023) [148] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
EarthPTi (2023) [149] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
USat (2023) [150] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
AIEarth (2023) [151] p ? $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Clay (2023) [152] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Hydro (2023) [153] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
U-Barn (2023) [154] $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
GeRSP (2023) [155] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
SwiMDiff (2023) [156] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
OFA-Net (2023) [157] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
SML-FR (2023) [158] p p $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Spectral-GP (2024) [18] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
S2MAE (2024) [159] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
SatMAE++ (2024) [160] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
msGFM (2024) [161] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
SkySense (2024) [24] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
MTP (2024) [162] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
DOFA (2024) [163] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
MMEarth (2024) [164] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
LeMeViT (2024) [165] $\times$ .. $\times$ $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
SoftCon (2024) [166] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
RS-DFM (2024) [167] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
A2MAE (2024) [168] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
HyperSIGM (2024) [169] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
SelectiveMAE (2024) [170] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
OmniSat (2024) [171] $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
MM-VSE (2024) [172] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
MA3E (2024) [173] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
SpectralEarth (2024) [174] $\times$ .. $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
SenPa-MAE (2024) [175] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
RingMo-Ae (2024) [176] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
SAR-JEPA (2024) [177] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
PIS (2024) [178] $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$ .. .. .. .. .. ..
OReole-FM (2024) [179] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
PIEVIT (2024) [180] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
SatVisionTOA (2024) [181] $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
RS-vHeat (2024) [182] $\times$ .. ? ? ? ? $\checkmark$ ? $\checkmark$ ? $\checkmark$
Prithvi-EO-2.0 (2024) [110] $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
AnySat (2024) [183] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
WildSAT (2024) [184] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
SeaMo (2024) [185] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
FoMo (2025) [186] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
SatMamba (2025) [187] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Galileo (2025) [188] p p $\checkmark$ ? $\checkmark$ $\times$ ? ? $\checkmark$ $\checkmark$ $\checkmark$
SatDiFuser (2025)[189] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
RoMA (2025) [190] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Panopticon (2025) [191] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
FedSense (2025) [192] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
Copernicus-FM (2025) [193] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
HyperFree (2025) [194] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
DynamicVis (2025) [195] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
FlexiMo (2025) [196] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Google Satellite (2025)[39] $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SkySense++ (2025) [197] $\times$ $\checkmark$ $\checkmark$ p p $\checkmark$ $\times$ p $\checkmark$ $\checkmark$ $\checkmark$
RingMoE (2025) [198] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$